OpenStudy (anonymous):

Add the vectors a= (0,6) and b= (10,0) and find the magnitude and angle from the horizontal of the resultant vector. Please help!

OpenStudy (anonymous):

it will be: a+b=(0+10,6+0)=(10,6) magnitude=(6^2 +10^2)^.5=11.66

OpenStudy (anonymous):

How do I find the angle?

OpenStudy (anonymous):

round it to nearest 10's

OpenStudy (anonymous):

or ones

OpenStudy (anonymous):

angle can be from by arctan (6/10) = 31.0

OpenStudy (anonymous):

found*

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

coco show ur work

OpenStudy (anonymous):

 the angle is arc tan (6/10)